OpenStudy (marigirl):

write z=1 - √3 i in polar form I got my r = 2 and i got arg (theta) value as -pi/3

OpenStudy (marigirl):

the answer reads: (g) The modulus of 1 - √3 i is √12 + (√3)2 = 2 and its argument θ satisfies tan θ = - √3/1 = - √3, whence θ is equal to either 2π/3 or 5π/3. Since y=-√3 is negative, we have θ = 5π/3. The polar form of -3 + 3i is thus 2 (cos 5π/3 + i sin 5π/3). ....

OpenStudy (marigirl):

 Thats what i thought